Metlock Services, Inc. is trying to establish the standard labor cost of a typical brake repair. The following data have been collected from time and motion studies conducted over the past month.
Actual time spent on the brake repairs 3 hours
Hourly wage rate $13
Payroll taxes 10% of wage rate
Setup and downtime 12% of actual labor time
Cleanup and rest periods 28% of actual labor time
Fringe benefits 20% of wage rate
Determine the standard direct labor hours per brake repairs. (Round answer to 2 decimal places, e.g. 1.25.)
Standard direct labor hours per brake repair

hours
Determine the standard direct labor hourly rate. (Round answer to 2 decimal places, e.g. 1.25.)
Standard direct labor hourly rate $

Determine the standard direct labor cost per brake repair. (Round answer to 2 decimal places, e.g. 1.25.)
Standard direct labor cost per brake repair $

If a brake repair took 3.60 hours at the standard hourly rate, what was the direct labor quantity variance? (Round answer to 2 decimal places, e.g. 1.25.)
Direct labor quantity variance $
Answer #1

1) Standard direct labor hours per brake repair = 3 Hour+(3*12%+3*28%) = 4.20 Hour

2) Standard direct labor hourly rate = 13+(13*30%) = $16.90

3) Standard direct labor cost per brake repair = 16.90*4.20 = $70.98

4) Direct labor quantity variance = (4.20-3.60)*16.90 = $10.14 F